Wendy has a special pair of dice. One each die 1 face is labeled head and 2 faces are labeled tail & the other three faces are labeled arm. If Wendy tosses the pair of dice, what is the probabilities she will get exactly 1 head and 1 arm?
If Wendy tosses her special pair of dice what is the probability that she will see at least 1 tail in the results?
1
Expert's answer
2011-11-18T08:43:14-0500
P(head)=1/6
P(tail)=1/3
P(arm)=1/2

P(x=head,y=arm)= P(x=head)* P(y=arm)=1/6*1/2=1/12
P(x=head,y=arm)= P(x=arm, y= head)=1/12
P=1/12+1/12=1/6
P(one tail)=P(x=tail)+P(y=tail)-P(x=tail,y=tail)=1/3+1/3-1/9=5/9